### v1

# UGM2N: An Unsupervised and Generalizable Mesh Movement Network via M-Uniform Loss

Abstract
[…]
1 Introduction
Solving partial differential equations (PDEs) is fundamental for modeling physical phenomena, spanning fluid dynamics, heat transfer, quantum mechanics, and financial markets [1]. Modern PDE solving critically relies on meshes, which serve as the foundational discretization framework for numerical methods [2, 3]. The accuracy and computational cost of PDE solutions are significantly affected by mesh resolution: high-resolution meshes resolve complex physics at high computational expense, whereas coarse meshes improve efficiency but risk missing critical features. As problem complexity grows, geometric details demand exponentially finer resolution, while multi-physics interactions require dynamic adaptation—pushing memory, parallel efficiency, and solver convergence to their limits [4]. To alleviate this issue, mesh adaptation methods—particularly mesh refinement method ($h$-adaptation method) and mesh movement method ($r$-adaptation method) —dynamically optimize computational resources, systematically overcoming traditional bottlenecks through intelligent spatial discretization control [5, 6, 7].
Mesh refinement method dynamically adjusts resolution via local element subdivision/coarsening, altering node counts while retaining fixed positions. In contrast, mesh movement method preserves node counts but relocates them strategically to high-resolution regions, guided by error estimators or gradients [8]. While mesh refinement method handles discontinuities via topological changes, mesh movement method suits smooth domains, avoiding remeshing overhead. However, the traditional Monge-Ampère (MA)-based methods suffer from high computational costs due to (1) repeated mesh-motion PDE solves (e.g., solving auxiliary equations) and (2) mesh-quality checks to prevent inversion. In extreme cases, adaptive operations can exceed the PDE-solving cost itself, making the enhancement of mesh adaptation efficiency an enduring open problem.
To improve the efficiency of mesh movement method, pioneering works employ supervised learning, training models on meshes adapted via traditional MA-based methods. Song et al. [9] propose a mesh adaptation framework trained via MSE loss between initial and adapted mesh nodes, and Zhang et al. [10] introduce a zero-shot adaptive model trained with a combined loss of volume preservation and Chamfer distance between initial and adapted mesh nodes. However, such supervised methods exhibit limited generalization: M2N requires PDE- and geometry-specific retraining, risking mesh tangling under extreme deformations, while UM2N’s zero-shot performance may degrade for unseen domains or PDEs.
In this paper, we propose UGM2N, an unsupervised and generalizable mesh movement network. Inspired by vision Transformers [11], we introduce node patches, locally normalized nodes with first-order neighbors, as model inputs. Unlike M2N/UM2N’s whole-mesh processing, our method parallelly and independently computes adapted positions for each patch, simplifying the learning objective and enabling scale-invariant mesh adaptation. Leveraging the node patch representation, we formulate an M-Uniform loss function that mathematically encodes local equidistribution properties, the core objective of mesh movement methods. Minimizing this patch-wise loss can produce approximately M-Uniform meshes while effectively matching MA-based adaptation objectives—all achieved through an efficient unsupervised framework. By learning adaptation dynamics directly, our model achieves equation-agnostic generalization while maintaining mesh-geometric independence.
Our main contributions are summarized as follows:
  • We present an unsupervised mesh movement network, eliminating the need for pre-adapted meshes by learning solely on initial meshes and flow fields. Our novel node-patch representation processes localized neighborhoods rather than full meshes, enabling efficient training and inherent generalization.
  • We derive a theoretically grounded M-Uniform loss function that enforces local mesh equidistribution at the node-patch level, which aligns with MA-based optimization objectives through a fully data-driven approach with native equation-agnostic generalization across arbitrary mesh geometry.
  • We present extensive numerical validation showing exceptional generalizability and robustness across various PDE types (both steady-state and time-dependent), accommodating different boundary conditions or initial conditions, and mesh geometries with varying shapes or resolutions.
2 Related Work
Machine learning for mesh generation and optimization. The automation and intelligence of mesh generation are among the key challenges in CFD 2030 [12], driving significant research efforts toward intelligent mesh generation and optimization. Zhang et al. [13, 14] proposed the MeshingNet and MeshingNet3D models to generate high-quality tetrahedral meshes, demonstrated in linear elasticity problems on complex 3D geometries. Chen et al. [15] introduced the MGNet model, which employs physics-informed neural networks [16] to achieve structured mesh generation. For mesh optimization, Guo et al. [17], Wang et al. [18, 19] developed intelligent mesh optimization agents based on supervised learning, unsupervised learning, and reinforcement learning (RL), achieving a balance between optimization efficiency and quality.
Machine learning for mesh adaptation. Unlike static mesh optimization methods, mesh adaptation dynamically modifies the computational mesh during simulation to enhance resolution in critical regions (e.g., shock waves, boundary layers, or vortex-dominated flows). These techniques are guided by error estimation schemes or feature-based criteria, ensuring computational efficiency while preserving accuracy. Advanced implementations leverage machine learning to predict optimal adaptation strategies, enabling high-fidelity simulations for complex, evolving flows.
For mesh refinement method, Foucart et al. [20] pioneered RL for adaptive mesh refinement, formulating it as a POMDP and training policy networks directly from simulations. Dzanic et al. [21] developed DynAMO, using multi-agent RL to predict future solution states for anticipatory refinement. Kim et al. [22] introduced GMR-Net, leveraging graph CNNs to predict optimal mesh densities without costly error estimation. Beyond these foundational works, research in intelligent $h$-adaptive mesh refinement remains highly active, with additional advancements documented in [23, 24, 25, 26, 27, 28, 29].
For mesh movement method, Omella and Pardo [30] proposed a neural network-enhanced boundary node optimization method, which is specifically designed for tensor product meshes. Song et al. [9] introduced M2N, a framework combining neural splines with GAT, enabling end-to-end mesh movement with 3–4 order-of-magnitude speedups. Hu et al. [31] introduced a neural mesh adapter trained via the MA equation physical loss to dynamically adjust mesh nodes, and develops a moving mesh neural PDE solver that improves modeling accuracy for dynamic systems. Rowbottom et al. [32] proposed a graph neural diffusion method that directly minimize finite element error to achieve efficient mesh adaptation. For specialized applications, methods such as Flow2Mesh and Para2Mesh have demonstrated the efficacy of learning-based adaptation in aerodynamic simulations [33, 34]. Recent work extended these advances with UM2N [10], a universal graph-transformer architecture attempts to achieve zero-shot adaptation across diverse PDEs and geometries. Most of the aforementioned works rely on supervised learning, where models are trained to align their outputs with pre-adapted meshes, resulting in a lack of physical information. Additionally, they often require retraining for different PDEs or mesh geometries, limiting their generalizability. This paper adopts an unsupervised learning approach to achieve equation-agnostic generalization across arbitrary mesh geometries.
3 Method
3.1 Problem statement and preliminaries
Given an initial mesh $\mathcal{M}$ (e.g., a uniform mesh) and associated flow field variables (such as velocity $\mathbf{u}$ or pressure $p$), the mesh movement method optimizes the node positions to generate an adapted mesh satisfying predefined resolution criteria. The mesh movement process can be analyzed from different perspectives, such as coordinate mapping between uniform and adapted meshes, uniform mesh construction in metric space, and so on [6]. The MA-based method adopts the former approach, solving the MA equation with boundary conditions to obtain the coordinate mapping. In contrast, this study employs the latter perspective, enforcing uniform distribution in metric space without explicit coordinate transformations between computational and physical domains.
From the latter perspective, given a physical domain $\Omega\subset\mathbb{R}^{d}$ (where $d\geq 1$), the goal of mesh movement is to construct uniform meshes in some metric space, which is defined by a matrix-valued monitor function $M=M(\mathbf{x})$, where $\mathbf{x}\in\Omega$. A mesh is said to satisfy the mesh equidistribution condition if it is uniformly distributed in this metric space, which can be mathematically expressed as:
$\int_{K}m(\mathbf{x})d\mathbf{x}=\frac{\sigma}{N_{e}},\forall K\in\mathcal{M},$ (1)
where $\sigma=\int_{\Omega}\rho(\mathbf{x})\,d\mathbf{x}$, $m(\mathbf{x})=\sqrt{\det(M(\mathbf{x}))}$ is the mesh density function, $K$ is the element of $\mathcal{M}$, and $N_{e}$ is the number of elements in the mesh $\mathcal{M}$. This condition constrains the size of mesh elements—when $m(\mathbf{x})$ is large, the element volume should be small, and vice versa. Additionally, the M-Uniform mesh condition also requires that the mesh elements should be equilateral in the metric space. In this work, we primarily focus on modifying the mesh density while disregarding the equilateral alignment of mesh elements.
Compared to MA-based coordinate transformations, this approach for constructing uniform meshes in the metric-space offers a more discretization-friendly framework, particularly well-suited for local loss function modeling (see Section 3.3).
3.2 Network overview
The proposed UGM2N is illustrated in Fig. 1. The model takes the initial mesh with solution as input, and node patches are constructed from all mesh nodes, with input features generated using flow field variables. The coordinates of mesh nodes within each patch are normalized to $0,1]\times[0,1]$ via 0-1 normalization, then encoded through node and edge encoders to obtain embeddings. These embeddings are processed by multiple deform blocks and a node decoder, producing adapted node coordinates for each patch, which are denormalized to restore the original mesh. The flow field features of the updated mesh are obtained through Delaunay Triangulation-based interpolation on the original mesh, and the adapted mesh serves as the initial input for the next iteration, with the process repeating for a maximum of $E$ epochs.
Figure 1: The proposed mesh movement network.
Node patches. Inspired by vision Transformers [11], the proposed model processes individual mesh node patches, unlike M2N or UM2N, which take the entire mesh as input. This patch-based approach significantly improves local feature representation while maintaining computational efficiency and scalability, mirroring the local optimization principles employed in mesh smoothing techniques.
In each adaptive epoch, the input consists of an initial mesh with a flow field solution, denoted as $\mathcal{M}=\{\mathcal{V},\mathbf{U},\mathcal{E}\}$, where $\mathcal{V}=\{\mathbf{x}_{1},\dots,\mathbf{x}_{N}\}$ represents node coordinates, $\mathcal{E}$ denotes connectivity, and $\mathbf{U}=\{\mathbf{u}_{1},\dots,\mathbf{u}_{N}\}$ contains flow variables on the nodes. A node patch $\mathcal{P}_{i}=\{\mathbf{X}_{i},\mathcal{E}_{i}\}$ is defined as the node itself, its first-order neighbors, and their connections (excluding inter-neighbor connections). Patch normalization scales nodes to a unit square, reducing learning difficulty and accommodating varying mesh sizes. It is worth noting that normalization does not introduce additional computational overhead, as it can be efficiently implemented using vectorized operations.
The flow field features are incorporated into the patch features using a mesh density function derived from the Hessian matrix:
$M(\mathbf{x}_{i})$ $=\left(1+\alpha\frac{\|\mathbf{H}(u_{i})\|}{\max_{j}\|\mathbf{H}(u_{j})\|}\right)I,$ (2)
$m(\mathbf{x}_{i})$ $=1+\alpha\frac{\|\mathbf{H}(u_{i})\|}{\max_{j}\|\mathbf{H}(u_{j})\|},$ (3)
where $\alpha$ is a constant, $u_{i}=\|\mathbf{u}_{i}\|_{2}$, $I$ is the identity matrix, and $\|\mathbf{H}(u_{i})\|$ is the Frobenius norm of the Hessian. The scalar $m(\mathbf{x}_{i})$ is concatenated with node coordinates (yielding a 3D input for 2D meshes) as the model’s input.
Mesh coordinate computation based on Graph Transformer model. The adapted node coordinates are computed using a lightweight model. Node and edge features (central-to-neighbor coordinate vectors) are first encoded via dedicated MLP encoders, followed by $L$ deform blocks for graph feature extraction. We employ a residual-connected Graph Transformer [35] in each block, proving effective despite its simplicity. The adapted patch coordinates are then computed through the node MLP decoder, and the centering mesh node within each patch (the pink node in Fig. 1) is denormalized to the original mesh space through vectorized operations. For boundary nodes, we ignore them and do not perform adaptation, since the output coordinates are unlikely to lie precisely on the boundary.
Iterative mesh adaptation with dynamic termination. Inspired by iterative mesh smoothing, multi-epoch mesh adaptation is employed to progressively refine node distribution during inference [^1] Multi-epoch mesh adaptation is disabled during training to simplify the training process., the Hessian norm values are updated between epochs via Delaunay triangulation-based interpolation from original to adapted nodes, providing initialization for subsequent adaptations. While this process could theoretically continue indefinitely, convergence is not guaranteed and mesh validity may degrade. A fixed epoch count would limit optimization capability; instead, we propose a metric-based adaptive strategy that dynamically determines termination based on optimization progress. This approach will be detailed following the presentation of our unsupervised loss function.
3.3 M-Uniform loss
Unlike existing methods, which adopted supervised loss functions to align predicted meshes with reference meshes, our approach addresses two key challenges in practical applications: (1) the frequent unavailability of high-quality reference meshes, especially for multi-physics or geometrically complex problems, and (2) the poor zero-shot generalization to novel PDE types beyond the training distribution. These limitations motivate our development of an unsupervised adaptation method.
As introduced in the Section 3.1, enforcing the mesh equidistribution condition offers a novel approach. Eq. 1 requires that the integral of $m(\mathbf{x})$ over any mesh element $K$ be constant. However, since $m(\mathbf{x})$ is only known at mesh nodes, exact integration is infeasible. We thus relax the strict M-Uniform condition to an approximate M-uniform condition:
$\int_{K}m_{K}d\mathbf{x}=m_{K}|K|=\frac{\sigma_{h}}{N_{e}},\forall K\in\mathcal{M},$ (4)
$m_{K}=\sqrt{\det(M_{K})},M_{K}=\frac{1}{|K|}\int_{K}M(\mathbf{x})d\mathbf{x},$ (5)
where $|K|$ represents the volume of the mesh element $K$ and $\sigma_{h}$ is a constant. Here, $M_{K}$ is approximated via nodal averages: for a triangular element $K$ with nodes $K_{1},K_{2},K_{3}$, $M_{K}=\frac{1}{3}\sum_{j=1}^{3}M(\mathbf{x}_{K_{j}}^{\prime})$ (note that $M(\mathbf{x}^{\prime})$ require interpolation to obtain), where $\mathbf{x}^{\prime}_{i}$ is the adapted position of node $i$ output by the model. Together with Eq. 2 and 3, we can obtain $m_{K}=\frac{1}{3}\sum_{j=1}^{3}m(\mathbf{x}_{K_{j}}^{\prime})$. Building upon these foundations, we can define a metric function for element $K$:
$\mathcal{L}_{K}=m_{K}|K|.$ (6)
Let $K_{l}^{i}$ be the mesh element $l$ in the patch of mesh node $i$. Rewriting Eq. 4 in terms of the local mesh node $i$, we require that $\mathcal{L}_{K^{i}_{l}}$ be as uniform as possible around mesh node $i$. We measure the variation in $\mathcal{L}_{K^{i}_{l}}$ among different mesh elements using a variance-based loss function:
$\mathcal{L}_{\mathrm{var}}\left(\mathcal{P}_{i}\right)=\frac{1}{N_{i}}\sum_{l=1}^{N_{i}}\left(\mathcal{L}_{K^{i}_{l}}-\overline{\mathcal{L}_{K^{i}_{l}}}\right)^{2},$ (7)
where $N_{i}$ is the number of mesh elements in the patch $\mathcal{P}_{i}$, and $\overline{\mathcal{L}_{K^{i}_{l}}}=\frac{1}{N_{i}}\sum_{l=1}^{N_{i}}\mathcal{L}_{K^{i}_{l}}$. Then, the proposed M-Uniform loss function can be writen as:
$\mathcal{L}_{M}\left(\mathbf{\theta}\right)=\lambda\mathbb{E}_{i\in\{1,\ldots,N\}}\mathcal{L}_{\mathrm{var}}\left(\mathcal{P}_{i}\right),$ (8)
where $\mathbf{\theta}$ is the model parameters, and $\lambda=100$ is a scaling constant. This approach shares conceptual similarities with PINNs, where local constraints—residual conditions in PINNs and the M-Uniform condition here—guide the learning process. By enforcing mesh equidistribution condition at the node level, the model can adapt mesh node positions without supervised data, ensuring generalization to arbitrary adaptive scenarios. Crucially, unlike existing adaptive methods (e.g., M2N or UM2N), training does not require the full mesh as input. Instead, it can be trained on individual mesh nodes. For example, during mini-batch training, we can sample a subset of mesh nodes from the mesh and achieve efficient training through graph batching. Moreover, this approach further reduces the required amount of data, as the number of data samples is proportional to the number of mesh nodes rather than the number of meshes.
During iterative mesh adaptation, we compute the global uniformity metric $\mathcal{L}_{\mathrm{var}}\left(\mathcal{M^{\prime}}\right)$ over the entire adapted mesh $\mathcal{M^{\prime}}$ after each epoch to assess equidistribution compliance:
$\mathcal{L}_{\mathrm{var}}\left(\mathcal{M^{\prime}}\right)=\frac{1}{N_{e}}\sum_{l=1}^{N_{e}}\left(\mathcal{L}_{K_{l}}-\overline{\mathcal{L}_{K_{l}}}\right)^{2},$ (9)
where $N_{e}$ is the number of mesh elements in $\mathcal{M^{\prime}}$, and $\overline{\mathcal{L}_{K_{l}}}=\frac{1}{N_{e}}\sum_{l=1}^{N_{e}}\mathcal{L}_{K_{l}}$. The model stops the iterative mesh adaptation when $\mathcal{L}_{\mathrm{var}}\left(\mathcal{M^{\prime}}\right)$ no longer decreases. During inference, we set a fixed upper limit for the number of iterations—specifically, we set the maximum number of mesh adaptation epochs as 10. The full adaptation algorithm is detailed in App. A.1, and the theoretical analysis of the effectiveness of the loss function to optimize mesh distribution is provided in App. A.2.
4 Experiment
4.1 Experiment setups
Different numerical simulations involve diverse flow-field and mesh geometries characteristics. An effective mesh movement method must account for variations in both the flow field and the underlying mesh geometry. Our experiments show that the proposed method achieves robust, optimal performance across both scenarios—whether applied to different flow fields (diverse PDEs with varying solutions) or entirely distinct mesh geometries.
Model training. Following UM2N’s protocol, we trained UGM2N on a mesh with only four flow fields and evaluated its zero-shot generalization performance on unseen flow fields or meshes. As depicted in Fig. 10 (App. B.1), random perturbations were applied to mesh node positions to enhance data diversity, resulting in a training set comprising 10,440 mesh nodes. The model was optimized using Nadam [36] with an initial learning rate of 1e-4, with all experiments conducted on an NVIDIA RTX TITAN GPU. See App. C for more training details.
Baselines and metrics. We performed a comparative analysis against the MA method [37], M2N, and UM2N, using UM2N’s pre-trained weights obtained from its GitHub repository [38]. Performance was evaluated using two key metrics: (1) error reduction (ER), which measures the relative improvement in PDE solution accuracy compared to the initial coarse mesh (with the high-resolution solution serving as the reference), and (2) tangling ratio (TR), defined as the fraction of invalid elements in the adapted mesh. Additional case-specific metrics will be presented in the corresponding experimental sections. Detailed mathematical definitions of all metrics are provided in App. F.
4.2 Performances across different flow field solutions
To assess the model’s generalization ability across diverse flow fields, we conducted experiments using Burgers’ equation with varying initial conditions, as well as Poisson and Helmholtz equations with different analytical solutions (refer to App. B.2 and B.3 for more detailed PDE configurations). All simulations were performed on a uniform triangular mesh spanning the domain $0,1]\times[0,1]$, comprising 1,478 elements. For validation, a high-fidelity reference solution was computed on a significantly refined mesh with 23,250 elements, ensuring precise accuracy for benchmarking purposes.
The test results are summarized in Table 1, it can be observed that the our model demonstrates significant advantages in the vast majority of cases. In the seven test cases for the Poisson equation, our model achieved optimal performance (highest ER or lowest TR) in five cases, particularly excelling with complex functions. For example, for $\sum_{i,j}\exp[-\frac{(x-x_{\mu,i})^{2}}{0.25^{2}}-\frac{(y-y_{\mu,j})^{2}}{0.25^{2}}]$, ours achieved an ER of 9% , far surpassing the comparison methods (MA: -8.64%, M2N: -30.20%).
Figure 2: Mesh adaptation results for the Helmholtz equation with different solutions.
Additionally, ours achieved complete dominance in the Helmholtz equation, securing the four highest ER out of five test cases. Although slightly inferior to M2N in the Burgers equation, ours still significantly outperformed UM2N. These results validate the robustness and generalization capability of our method in mesh adaptation across different PDEs, achieving state-of-the-art performance compared to existing approaches.
The mesh adaptation results for the Helmholtz equation are shown in Fig. 2 (for results on the Poisson variance and Burgers equation, refer to App. D). Our method demonstrates superior visual alignment with the target flow field while simultaneously generating meshes with excellent quality compared to alternative approaches. It is noteworthy that although M2N and UM2N can also produce approximately adaptive results, they show weaker adherence to the mesh equidistribution condition compared to the our method (see App. E).
Table 1: Model performance of different flow fields
[…]
MA [37] M2N [9] UM2N [39] Ours
$u_{\text{exact}}$
Poisson $1+8\pi^{2}\cos(2\pi x)\cos(2\pi y)$ 15.40 0.92 6.74 14.56
[…]
$\sin(4\pi x)\sin(4\pi y)$ 9.79 -98.01 -2.19 12.46
$1/\exp((x-0.5)^{2}+(y-0.5)^{2})$ -28.22 1.15 -2.98 1.70
$\sin(2\pi x+2\pi y)$ 11.69 -34.03 -9.07 9.07
$\cos(\pi x)\exp\!\left(-((x-0.5)^{2}+(y-0.5)^{2})\right)$ -8.94 -41.89 5.15 4.90
$\cos\!\left(\sqrt{(x-0.5)^{2}+(y-0.5)^{2}}\right)\times$
$\exp\!\left(-((x-0.5)^{2}+(y-0.5)^{2})\right)$
-25.23 1.62 -3.53 2.56
$u_{\text{exact}}$
Helmholtz
$\cos(2\pi y)$ 15.60 -11.16 10.86 14.11
$\cos(2\pi x)$ 10.29 -37.24 6.80 13.15
$\cos(2\pi y)\cos(2\pi x)$ 13.48 -24.33 5.63 15.03
$\cos(2\pi y)\cos(4\pi x)$ 10.87 -351.63 -2.61 14.09
$\cos(4\pi y)\cos(2\pi x)$ 13.50 -250 3.43 16.98
$\mathbf{u}_{\text{ic}}$
Burgers
$\left[\sin\left(-20\left(x-0.5\right)^{2}\right),\cos\left(-20\left(y-0.5\right)^{2}\right)\right]^{T}$ 26.81 44.82 0.46 32.22
$\left[\exp\left(-\left(\left(x-0.5\right)^{2}+\left(y-0.5\right)^{2}\right)\times 100\right),0\right]^{T}$ 51.12 29.93 22.76 30.19
4.3 Performance across varying mesh geometries
Figure 3: The ER for different mesh resolutions on the Helmholtz equation. For clarity in presentation, we clipped the minimum ER at -1%, even though for some methods (e.g., M2N), their adapted meshes significantly increased the solution error.
Figure 4: The performance of mesh movement methods under different mesh resolutions is presented. The dashed line in the figure indicates the time required for UGM2N to complete a single iteration.
Varying mesh resolutions. Practical simulations often involves meshes with varying element sizes, making it crucial for the model to handle meshes of different resolutions. We tested the model’s performance on the Helmholtz equation with different mesh element sizes, where the solutions are the same as in Table 1. The coarse mesh element sizes were [0.05, 0.04, 0.03, 0.02], corresponding to mesh element counts of [944, 1478, 2744, 5824]. As shown in Fig. 4, our model achieved improved solution accuracy across all element sizes, whereas the M2N model failed to adapt the mesh at any resolution, and UM2N could only generalize on some of the element sizes. The results demonstrate that the loss function based on MSE between mesh nodes in M2N struggles to generalize to unseen meshes. Furthermore, UM2N’s volume loss exhibits only limited generalizability.
Regarding computational efficiency, as illustrated in Fig. 4, we present the time required for mesh movement under varying element sizes. For each model configuration, we conducted ten repeated tests on different solutions of the Helmholtz equation and reported the average mesh movement time per trial. Compared with the MA method, the neural-based mesh movement method demonstrates significant advantages. Although the iterative mesh adaptation in UGM2N introduces computational overhead proportional to the optimization epoch $E$, its computational efficiency does not exhibit substantial degradation when compared to M2N or UM2N. The reason lies in its relatively low per-epoch optimization time (indicated by the dashed line in the figure).
Varying mesh shapes. Another potential variable in the simulation is the shape of mesh. To evaluate the model’s generalization capability under different mesh geometric configurations, we quantitatively conducted three distinct simulation cases: subsonic flow around a NACA0012 airfoil, cylinder flow, and the wave equation on a circular domain, with the experimental setups provided in App. B.2. The adaptive results of the airfoil mesh are shown in Fig. 6, which shows that our model effectively captures the shock wave location without introducing any invalid elements. Additionally, at the position $y=0.25$, our model obtains the minimum mean absolute error (MAE  [^3] The average absolute error between the solution on the high-resolution mesh and the solution on the adaptive (coarse) mesh at each position (time step).) in pressure coefficient result. For the cylinder flow case, neither MA nor M2N could produce valid meshes. As shown in Fig. 6, compared to UM2N, our model further reduces the prediction error of the drag coefficient and slightly decreasing the cumulative error during the solving process. The results of the wave equation are shown in Fig. 8. At any given time step, our proposed method consistently generates smooth, high-quality adaptive meshes. In contrast, other methods may produce distorted mesh elements and fail to reduce errors, although they can also generate adaptive meshes.
Figure 5: Results on the cylinder flow. We present the adaptive meshes at four time slices, with the results over the entire simulation period provided in the supplementary vedio materials.
Figure 6: Mesh adaptation on the subsonic flow case. MA method fails to converge in this case.
Figure 7: Mesh adaptation results on the wave equation.
Figure 8: Mesh adaptation on moving cylinder and supersonic flow cases. In supersonic flow, UGM2N’s adaptive mesh dynamically adjusts nodes to track shock waves; in the moving cylinder case, it precisely captures both the cylinder boundary and wake flow. The results are also presented in the supplementary vedio.
In addition to the above quantitative experiments, two additional qualitative experiments—supersonic flow over a wedge and a moving cylinder inside a channel—are provided in Fig. 8; the results also demonstrate that our model can also generate effective adaptive meshes for more complex cases.
4.4 Ablation Study
Loss function To validate the effectiveness of the M-Uniform loss compared to the coordinate loss of M2N and the volume loss of UM2N, we trained models with the same architecture and the same training data but different loss functions. The supervised data was generated using the MA method. Table 9 presents the average ER on PDEs with different solutions of models trained with different loss functions on three types of equations (the equation configurations are the same as in Table 1). With only a small amount of training data, supervised learning methods using the entire mesh as input struggle to produce effective models. In contrast, our unsupervised training approach requires no adaptive meshes as supervised data, and the node patch-based training method enables effective model training even with limited data.
Iterative mesh adaptation To demonstrate the effectiveness of our iterative mesh adaptation approach, Fig. 9 shows the error reduction across optimization iterations for the Poisson equation in Table 1. The results reveal that the error reduction exhibits a non-monotonic trend, initially increasing before decreasing as optimization progresses. Notably, our adaptive adaptation epochs (marked with a diamond) consistently stay within high ER regime, demonstrating the effectiveness of our method.
Table 2: The mesh adaptation performance of models trained with different loss functions
ER (%) $\uparrow$
Loss Poisson Helmholtz Burgers
Coordinate loss -8.19 -4.46 -9.17
Volume loss -8.27 -0.52 -1.46
M-Uniform loss 5.21 9.94 30.07
Figure 9: Error reduction in solving the Poisson equation under different adaptive epoch settings.
5 Conclusion
We introduce UGM2N, an unsupervised and generalizable mesh movement network. This network removes the requirement for pre-adapted meshes while demonstrating strong generalization capabilities across various PDEs, geometric configurations, and mesh resolutions. By leveraging localized node-patch representations and a novel M-Uniform loss, our approach enforces mesh equidistribution properties comparable to Monge-Ampère-based methods—but in a more efficient, unsupervised manner. Extensive experiments demonstrate consistent performance improvements over both supervised learning baselines and traditional mesh adaptation techniques, achieving significant error reductions without mesh tangling across diverse PDEs and mesh geometries. See App. G for the discussions on limitations and broader impacts.
Appendix A Method
A.1 The proposed mesh movement method
The proposed mesh movement method based on UGM2N is presented in Alg. 1. Given an initial mesh and the corresponding flow field, UGM2N iteratively refines the mesh through multiple adaptive operations to obtain the optimal mesh. In this algorithm, operations such as Patch Processing, Mesh Reconstruction, and Convergence Check can all be vectorized, ensuring high computational efficiency in our method.
Algorithm 1 The proposed mesh movement method
0: Initial mesh $\mathcal{M}^{0}=\{\mathcal{V}^{0},\mathbf{U}^{0},\mathcal{E}\}$, max epochs $E$
0: Adapted mesh $\mathcal{M}^{*}$
1:for $e=1$ to $E$ do
2:  Patch Processing:
3:  Construct node patches $\{\mathcal{P}_{i}\}_{i=1}^{N}$ from $\mathcal{M}^{e-1}$
4:  for each patch $\mathcal{P}_{i}=\{\mathbf{X}_{i},\mathcal{E}_{i}\}$  do
5:   Normalize coordinates $\mathbf{X}_{i}\rightarrow[0,1]^{d}$ // Vectorized operations in the loop
6:   Compute density function $m(\mathbf{x}_{i})$ via Eq. 2 and 3
7:   Encode features: $\mathbf{H}_{i}=\text{NodeEncoder}(\left[\mathbf{X}_{i},m(\mathbf{X}_{i})\right])$ // The operator $m$ is applied row-wise to the $\mathbf{X}_{i}$
8:   Update positions: $\mathbf{X}^{\prime}_{i}=\text{NodeDecoder}(\text{DeformBlocks}(\mathbf{H}_{i},\text{EdgeEncoder}(\mathcal{E}_{i})))$
9:   Denormalize centering node in $\mathbf{X}^{\prime}_{i}$ to the original mesh space
10:  end for
11:  Mesh Reconstruction:
12:  Assemble adapted mesh $\mathcal{M}^{\prime}=\{\mathcal{V}^{\prime},\mathbf{U}^{\prime},\mathcal{E}^{\prime}\}$
13:  Interpolate Hessian norm (or grad norm) $\|\mathbf{H}(\mathbf{x}^{\prime}_{i})\|$ via Delaunay triangulation
14:  Convergence Check:
15:  Compute global uniformity $\mathcal{L}_{\mathrm{var}}(\mathcal{M}^{\prime})$ via Eq. 9
16:  if $\mathcal{L}_{\mathrm{var}}$ stops decreasing or $e==E$ then
17:   $\mathcal{M}^{*}\leftarrow\mathcal{M}^{\prime}$
18:   break
19:  else
20:   $\mathcal{M}^{e}\leftarrow\mathcal{M}^{\prime}$
21:  end if
22:end for
A.2 Analysis of M-Uniform loss
Here, we theoretically demonstrate that optimizing the local M-Uniform loss effectively optimizes the objective function associated with mesh equidistribution in one adaptation epoch. Consider a mesh $\mathcal{M}$ (we omit epoch $e$ for clarity), let $L_{K}\in\{L_{K^{i}_{j}}\mid i\in\{1,2,\dots,N\},j\in\{1,2,\dots,N_{i}\}\}$ and $L_{K}^{\prime}\in\{L_{K_{l}}\mid l\in\{1,2,\dots,N_{e}\}\}$ represent discrete random variables defined over mesh elements (see Eq. 6 and Eq. 9), and $I\in\{1,2,\dots,N\}$ be a discrete uniform random variable. Here, $N$ denotes the total number of mesh nodes, while $N_{i}$ indicates the number of mesh elements in the patch centered at node $i$. Simply put, $L_{K}^{\prime}$ is a random variable defined on all mesh elements, and $L_{K^{i}_{j}}$ is a random variable defined on the mesh elements contained within all patches. According to the law of total variance, we have:
$\operatorname{Var}(L_{K})=\mathbb{E}[\operatorname{Var}(L_{K}\mid I)]+\operatorname{Var}\left(\mathbb{E}[L_{K}\mid I]\right).$ (10)
The expectation $\mathbb{E}[\operatorname{Var}(L_{K}\mid I)]$ quantifies the average local variance across node-centered patches, which simplifies to:
$\mathbb{E}\left[\operatorname{Var}\left(L_{K}\mid I\right)\right]=\frac{1}{N}\sum_{i=1}^{N}\left(\underbrace{\frac{1}{N_{i}}\sum_{l=1}^{N_{i}}(L_{K_{l}^{i}}-\overline{{L}_{K_{l}^{i}}})^{2}}_{\mathcal{L}_{\mathrm{var}}(P_{i})}\right)=\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}_{\mathrm{var}}(P_{i}).$ (11)
Moreover, from Eq. 9, we have $\mathcal{L}_{\mathrm{var}}(\mathcal{M})=\operatorname{Var}({L_{K}^{\prime}})$. When the samples in $L_{K}$ are repeated samples from $L_{K}^{\prime}$ (i.e., each sample is duplicated $n$ times), we have $\operatorname{Var}(L_{K})=\operatorname{Var}(L_{K}^{\prime})$. Assuming that each mesh element appears in approximately the same number of local patches—i.e., the sampling of $L_{K}^{\prime}$ in $L_{K}$ is nearly identical—we can adopt the approximation $\operatorname{Var}(L_{K})\approx\operatorname{Var}(L_{K}^{\prime})$. This assumption holds, for example, in high-quality triangular meshes generated by mesh generation software, where almost all mesh nodes have a degree of 6. Moreover, the number of nodes on the boundary is relatively small compared to the number of interior nodes. With $\operatorname{Var}\left(\mathbb{E}[L_{K}\mid I]\right)\geq 0$, we get such an inequality:
$\mathcal{L}_{\mathrm{var}}(\mathcal{M})=\operatorname{Var}(L_{K}^{\prime})\approx\operatorname{Var}(L_{K})\geq\mathbb{E}[\operatorname{Var}(L_{K}\mid I)]=\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}_{\mathrm{var}}(P_{i})=\frac{1}{\lambda}\mathcal{L}_{M}\left(\mathbf{\theta}\right),$ (12)
$\lambda\mathcal{L}_{\mathrm{var}}(\mathcal{M})\geq\mathcal{L}_{M}\left(\mathbf{\theta}\right).$ (13)
Therefore, $\mathcal{L}_{M}(\theta)$ provides a valid lower bound for $\lambda\mathcal{L}_{\mathrm{var}}(\mathcal{M})$. In each adaptation epoch, when the model can successfully minimizes $\mathcal{L}_{M}(\theta)$, it concurrently optimizes the lower bound of $\lambda\mathcal{L}_{\mathrm{var}}(\mathcal{M})$, thereby promoting mesh equidistribution. Moreover, in the limit where $\mathcal{L}_{M}(\theta)=0$, all local variances $\mathcal{L}_{\mathrm{var}}(P_{i})$ vanish, implying that $L_{K}$ is constant over all patches. Consequently, $L_{K}^{\prime}$ must also be constant, leading to $\lambda\mathcal{L}_{\mathrm{var}}(\mathcal{M})=0$, which corresponds to exact mesh equidistribution.
Appendix B Dataset
B.1 Train data setups
We only used four flow fields on the same mesh to train the model, as shown in Fig. 10. The analytical solutions for the four flow fields are:
  1. 1.
    $u_{1}(x,y)=10\sin(2\pi x)\sin(2\pi y)$;
  2. 2.
    $u_{2}(x,y)=-\frac{5}{\pi}\sin(\pi x)\sin(\pi y)$;
  3. 3.
    $u_{3}(x,y)=10(\sin(5x)^{10}+\cos(10+25xy)\cos(5x))$;
  4. 4.
    $u_{4}(x,y)=10(1-e^{x}\cos(4\pi y))$
Additionally, we perform data augmentation on the mesh by introducing random perturbations to the mesh nodes.
Figure 10: Flow flieds for trianing the mesh movement model.
B.2 Test data setups
Helmholtz. The Helmholtz equation describes the propagation of time-harmonic waves in physics and engineering. Here, we solve an equation of the following form:
$\nabla^{2}u+u=f,u=g\ \text{on}\ \partial\Omega,$ (14)
where $u$ is the solution variable, $\partial\Omega$ denotes the boundary, $g$ is the boundary function for $u$, and $f$ is the source term. To test the generalization capability of the model, we construct five different solutions (see Table 1), compute $f$ and $g$ to formulate the Helmholtz equation, and evaluate the model’s performance.
Poisson. The Poisson equation describes how a scalar field responds to a given source distribution, written as $\nabla^{2}u=f$. Using the same approach as for the Helmholtz equation, we constructed Poisson equations with seven different exact solutions (see Table 1) to test the model.
Burgers. The Burgers equation is a fundamental nonlinear partial differential equation in fluid dynamics, combining convection and diffusion effects. In this paper, we solve the following Burgers equation:
$\frac{\partial\mathbf{u}}{\partial t}+(\mathbf{u}\cdot\nabla)\mathbf{u}-\nu\nabla^{2}\mathbf{u}$ $=0,$ (15)
$(\mathbf{n}\cdot\nabla)\mathbf{u}$ $=0\text{ on }\Omega,$
where the viscosity coefficient $\nu=0.005$, and the initial conditions are given in Table 1. The simulation employs a time step of $\Delta t=\frac{1}{30}s$ and runs for a total duration of $0.5s$.
Airfoil and cylinder flow case. The airfoil case was simulated under conditions of Mach 0.8 at 1.55° angle of attack, while the cylinder flow case employed a Reynolds number of 100 with characteristic length of 0.2, kinematic viscosity of 0.01, time step of $0.001s$, and total simulation duration of $3.5s$. In our experiments, the airfoil case employed a coarse mesh with 10,466 elements and a high-resolution mesh with 41,364 elements, while the cylinder flow used 5,536 (coarse) and 11,624 (high-resolution) elements.
Wave equation. The wave equation is a partial differential equation that describes the propagation of waves (such as sound waves, light waves, water waves, etc.) through a medium or space. This experiment solves the two-dimensional wave equation for the initial value problem on a unit circle, namely:
$\begin{aligned} \frac{\partial^{2}u}{\partial t^{2}}-\nabla^{2}u=0\end{aligned},$ (16)
where the initial condition is $u(x,y,0)=(1-x^{2}-y^{2})\sin(\pi x)\sin(\pi y)$. The time step is $0.01s$, and the total time is $2s$. The coarse mesh consists of 2,048 elements, while the high-resolution mesh contains 8,192 elements.
Moving cylinder and supersonic flow over the wegde. Both cases are benchmark cases from the Kratos official website [40]. The mesh element count for the moving cylinder case is 9,852, and the element count for the supersonic flow case is 27,115. Please refer to the official site for more details.
B.3 Monitor function setups
Table 3 presents the values of the monitoring function $\alpha$ for different cases. Notably, in the airfoil case study, the monitoring function employs the gradient norm (rather than the Hessian norm) of mesh nodes to better capture the shock location. To address the long-tail distributions observed in both the cylinder flow and airfoil cases (since most mesh nodes have small Hessian norm values), a logarithmic transformation was applied to the monitoring function at mesh nodes.
Table 3: $\alpha$ for different cases
Case $\alpha$
Train data 5
Poisson, Helmholtz, Burgers, Wave 5
Cylinder flow, Airfoil case 10
Appendix C More model training details
We partitioned the dataset into training, validation, and test sets with an 8:1:1 ratio, employing the validation set for early stopping. The model architecture consists of node and edge encoders implemented as two linear layers [2, 512], followed by deformation blocks containing an 8-layer Graph Transformer with 512 hidden dimensions, residual connections, and 4 attention heads, and finally a node decoder comprising a LayerNorm-equipped MLP [512, 256, 2]. All components utilize ReLU activation functions. The model was trained on a machine with an I7-9700KF CPU, NVIDIA RTX TITAN GPU, and 64GB of memory, with the complete model only requiring approximately 3.1 hours of training time.
Appendix D Mesh adaptation results on Poisson’s equation and Burgers’ equation
As shown in Fig. 11, our method can effectively generate adaptive meshes in flow fields with different distributions, whether for steady or unsteady cases.
Figure 11: Mesh adaptation on Poisson’s equation and Burgers’ equation.
Appendix E The mesh equidistribution condition on the adapted mesh
Fig. 12 shows the absolute error between $\mathcal{L}_{K}$ and $\overline{\mathcal{L}_{K}}$ on each mesh element after adaptation, while Table 4 presents the value of $\mathcal{L}_{\mathrm{var}}\left(\mathcal{M^{\prime}}\right)$ on the mesh. Quantitative analysis reveals our method produces minimal error values, indicating its adapted mesh most closely approximates the ideal equidistribution condition.
Figure 12: $|\mathcal{L}_{K}-\overline{\mathcal{L}_{K}}|$ on the adapted meshes.
Table 4: $\mathcal{L}_{\mathrm{var}}\left(\mathcal{M^{\prime}}\right)$ on the adapted meshes for the Helmholtz equation with different solutions
Solution of the Helmholtz equation
Method $\cos(2\pi y)$ $\cos(2\pi x)$ $\cos(2\pi y)\cos(2\pi x)$ $\cos(2\pi y)\cos(4\pi x)$ $\cos(4\pi y)\cos(2\pi x)$
MA [37] 3.21e-7 3.89e-7 1.68e-7 1.30e-7 1.50e-7
M2N [9] 4.41e-7 5.90e-7 2.62e-7 4.67e-7 4.82e-7
UM2N [39] 5.41e-7 5.50e-7 2.89e-7 3.32e-7 3.73e-7
UGM2N 2.12e-7 2.68e-7 1.52e-7 1.10e-7 1.19e-7
Appendix F Evaluation metrics
Error reduction (ER). Error reduction quantifies the relative enhancement in PDE solution accuracy, computed as the improvement over the initial coarse mesh solution, where the high-resolution result is treated as the ground truth. For the steady case, error reduction is defined as:
$\text{ER}=\frac{\|\mathbf{u}_{\text{coarse }}-\mathbf{u}_{\text{ref }}\|_{2}-\left\|\mathbf{u}_{\text{adapted }}-\mathbf{u}_{\text{ref }}\right\|_{2}}{\left\|\mathbf{u}_{\text{coarse }}-\mathbf{u}_{\text{ref }}\right\|_{2}}\times 100\%,$ (17)
[…]
$\text{ER}=\frac{\sqrt{\sum_{i}^{T}\|\mathbf{u}_{\text{coarse},i}-\mathbf{u}_{\text{ref},i}\|^{2}_{2}}-\sqrt{\sum_{i}^{T}\left\|\mathbf{u}_{\text{adapted},i}-\mathbf{u}_{\text{ref},i}\right\|^{2}_{2}}}{\sqrt{\sum_{i}^{T}\|\mathbf{u}_{\text{coarse},i}-\mathbf{u}_{\text{ref},i}\|^{2}_{2}}}\times 100\%,$ (18)
where $\mathbf{u}_{\text{coarse},i}$, $\mathbf{u}_{\text{adapted},i}$, and $\mathbf{u}_{\text{ref},i}$ represent the numerical solutions at timestep $i$ from the initial coarse mesh, adapted mesh, and high-resolution mesh respectively.
Cumulative error. Cumulative error refers to the accumulation of errors over time, which is defined as:
$\text{Cumulative error}=\sqrt{\sum_{i}^{T}\|\mathbf{u}_{i}-\mathbf{u}_{\text{ref},i}\|^{2}_{2}},$ (19)
where $\mathbf{u}_{i}$ and $\mathbf{u}_{\text{ref},i}$ are the numerical solutions at timestep $i$ from the current mesh and the high-resolution mesh, and $T$ is the total timestep.
MAE of $C_{p}$ and $C_{D}$. The MAE (Mean Absolute Error) of the pressure coefficient $C_{p}$ and drag coefficient $C_{D}$ measures the errors between the solutions obtained at different positions or time steps and the results from the high-resolution mesh, defined as:
$C_{p,\text{MAE}}=\text{Mean}_{i}|C_{p,i}-C_{p,i}^{\text{ref }}|,C_{D,\text{MAE}}=\text{Mean}_{i}|C_{D,i}-C_{D,i}^{\text{ref }}|,$ (20)
where $C_{p,i}$ and $C_{D,i}$ represent the pressure coefficient and drag coefficient computed on the coarse (adapted) mesh at the $i$-th location (or time step), and $C_{p,i}^{\text{ref }}$ and $C_{D,i}^{\text{ref }}$ denote the corresponding values computed on the high-resolution mesh at the same location/time step.
Appendix G Limitations and broader impacts
Limitations. 1) The model does not handle mesh nodes on the boundary; future work should consider how to achieve the movement of boundary mesh nodes. 2) The approximation in Eq. 12 relies on the assumption that "each mesh elment appears approximately the same number of times in local patches." However, this condition only holds when the degrees of mesh nodes are nearly identical. Therefore, the method still requires further testing on non-uniform meshes (e.g., where mesh node degrees vary significantly). 3) The model adopts a relatively simple and lightweight architecture. Future work could explore more complex model designs to achieve better mesh adaptation performance.
Broader impacts. The UGM2N method proposed in this paper significantly improves the efficiency of mesh movement techniques in mesh adaptation through unsupervised learning and localized mesh adaptation technology. It reduces the high computational costs of traditional mesh adaptation methods, thereby accelerating the simulation of complex physical phenomena such as fluid dynamics and heat transfer under limited computational resources. Its generalizability allows it to be applied to various partial differential equations and geometric shapes, enhancing the practical engineering applications of current AI-based mesh adaptation methods.